Roscoe must play 24 commercials during his 1-hr radio show. Each commercial is either thirty seconds or sixty seconds long. If the total commercial time during that hour is 16 minutes, how many commercials of each type does Roscoe play?
Answer by mananth(16946)

x= commercials of 30 sec
y= commercials of 60 sec kid ticket
Total comercials
1 x + 1 y = 24 .............1
Total value
30 x + 60 y = 960 .............2
Eliminate y
multiply (1)by -60
Multiply (2) by 1
-60 x -60 y = -1440
30 x + 60 y = 960 ( 16 mins = 960 sec)
Add the two equations
-30 x = -480
/ -30
x = 16
plug value of x in (1)
1 x + 1 y = 24
16 + y = 24
y = 24 -16
y = 8
y = 8
x= 16 commercials of 30 sec
y= 8 commercials of 60 sec
m.ananth@hotmail.ca